OpenStudy (anonymous):

Suppose a boat travels due north for 17 miles, then heads northeast for 8 miles. How far from the original starting position is the boat? I know this deals with law of Cosines, but I'm very confused. I tried the distance formula, but I got 26 which I know is wrong as it can't be more than the other two sides added together which would be 25.

OpenStudy (campbell_st):

draw a diagram 

OpenStudy (anonymous):

Oh wow, thank you! I got uhm

OpenStudy (campbell_st):

so the distance will be the longest side of the triangle... because it is opposite the largest angle...

OpenStudy (anonymous):

Yes, I understand that. I'm a bit confused as to which formula to use to solve. Sorry we just got into law of Cosine and I'm horrible with trig!

OpenStudy (campbell_st):

well you have the measure of 2 sides and the included angle so its \[c^2 = a^2 + b^2 - 2ab \times \cos(C)\]

OpenStudy (anonymous):

64+289-2(8)(17)xCos(135) Right?

OpenStudy (anonymous):

623.935891158 Right?

OpenStudy (campbell_st):

that looks good... but don't forget thats c^2 so you need to square root

OpenStudy (anonymous):

Which I got uhm 24.9787087568

OpenStudy (campbell_st):

well its a sensible answer... as its the longest side... I haven't calculated it...

OpenStudy (anonymous):

Well, can you check for me?